Between z = 0 and z = −0.32

Between z = 0 and z = −0.32
Answer #1

Solution: It is required to find the area between z = 0 and z = -0.32

Z=-0.32 z=0

P(-0.32<z<0)=P(z<0)-P(z<-0.32)

Now using the standard normal table, we have:

\boldsymbol{P(-0.32<z<0)=P(z<0)-P(z<-0.32)=0.50-0.3745=0.1255 }

Therefore, the area between z = 0 and z = -0.32 is 0.1255